Thomas is 12 years older than Richard four years ago. Thomas was four times asold as Richard find their present ages
Answer by advanced_Learner(501)

richard age x
thomas age x+12
four year ago
x-4 richard age
x+12-4 thomas age
x+12-4=4(x-4)
x+8=(4x-16)
24=3x
x=8
check
richard age 8
thomas age 8+12 or 20
four year ago
x-4 ,8-4 or 4 is richard age
x+12-4 ,x+8 ,8+8 or 16 is thomas age
notice that thomas x+12-4 16 is indeed four times 4